# Exploring Imbalanced Annotations for Effective In-Context Learning

### v1

@@ 4 Methodology
In this section, we first decompose the distributional differences between annotated and test datasets to explain why classical rebalance methods fail for ICL. In light of this, we propose a novel method to enhance the performance of ICL with imbalanced annotated datasets.
@@ 4.1 Decomposition of the distributional differences
We aim to design a suitable demonstration selection method, which is used to select demonstrations for the given test input $\mathrm{\mathbf{x}}_{t}$, to minimize expected $\operatorname{Error}$ value across all classes:
where $M[\cdot,\cdot]$ measures the mismatched degree between real output and the generated output conditioned on demonstrations $\mathrm{\mathbf{C}}_{K}$ and test input $\mathrm{\mathbf{x}}_{t}$. For example, for classification tasks, $M[\cdot,\cdot]$ denotes error rate. For question answering tasks, $M[\cdot,\cdot]$ represents the negative value of EM score.
We apply the importance sampling trick to connect the expected $\operatorname{Error}$ and decompose the distributional differences into two-component weights as follows:
@@
$\mathbb{E}_{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right],$
@@
$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$
@@
where $P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})/P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})$ measures the distributional difference, $\bm{w}=P_{t}(Y)/P_{c}(Y)$ denotes class-wise weights and $\bm{\beta}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ is conditional bias. The detailed proof can be found in Appendix A.2.
Notably, these existing rebalance methods (i.e. re-weighting (Cui et al., 2019)) assume a shared conditional distribution between annotated and test datasets $P_{c}(X|Y)=P_{t}(X|Y)$ and only focus on how to use the class-wise weights $\bm{w}$ to process imbalanced datasets. However, in the real world, the conditional distribution $P(X|Y)$ is also prone to changes from the annotated dataset to the test dataset due to target shift and concept shift, especially in imbalanced datasets (Gopalan et al., 2013; Jamal et al., 2020; Gu et al., 2022).
We also conduct experiments to verify whether there is a shared conditional distribution between annotated and test datasets $P_{c}(X|Y)=P_{t}(X|Y)$ in imbalanced datasets. Specifically, for each class, we firstly transform text input from each class into bag-of-words representation, employ the Latent Dirichlet Allocation (LDA) model to derive the document-topic distributions and then compute Kullback-Leibler (KL) divergence between annotated conditional distribution $P_{c}(X|Y)$ and test conditional distribution $P_{t}(X|Y)$. For example, as shown in Figure LABEL:figure:4, the KL divergence between annotated and test datasets is unequal to zero and the assumption that $\beta=0$ is difficult to be satisfied.
In this paper, we relax the above assumption and allow a different conditional distribution of annotated and test datasets (e.g., $P_{c}(X|Y)\neq P_{t}(X|Y)$). However, directly computing the conditional bias $\bm{\beta}$ is non-trivial as real conditional distributions of annotated and test datasets are unknown. To circumvent the issue, we aim to design an effective method to estimate the conditional bias $\bm{\beta}$.
4.2 Modeling the conditional bias $\bm{\beta}$
Motivated by the previous analysis, we employ two-component weights, including class-wise weights $\bm{w}$ and conditional bias $\bm{\beta}$, to modify the original scoring functions during selection. We let the class-wise weights resemble the empirically successful design in the literature. Our key idea is to estimate conditional bias $\bm{\beta}$ by achieving empirical $\operatorname{Error}$ minimization of a balanced validation set.
Constructing balanced validation set
@@ 4.2 Modeling the conditional bias $\bm{\beta}$
Given an imbalanced dataset $\mathcal{D}_{c}$ with $N$ examples, the balanced validation set $\mathcal{D}_{b}$ is constructed by randomly selecting $n_{b}$ examples from $j$-class dataset $\mathcal{D}_{j}$ as follows:
$\mathcal{D}_{b}=\cup_{j=1}^{k}\left\{\operatorname{RandomSample}(\mathcal{D}_{j},n_{b})\right\},$ (7)
where $\operatorname{RandomSample}(\mathcal{D}_{j},n_{b})$ denotes randomly selecting $n_{b}$ examples from each class set $\mathcal{D}_{j}$. Notably, $n_{b}$ is constrained by the condition $n_{b}<\min(n_{1},...,n_{k})$. This process results in a validation set that is balanced across all classes, with each class contributing the same number of examples, similar to the composition of the test dataset. Meanwhile, $\mathcal{D}_{r}$ denotes the imbalanced remaining dataset.
@@
Imbalance Ratio 0 10 50 100 0 10 50 100
Random 32.63$\pm$0.73 32.26$\pm$1.32 30.91$\pm$1.20 30.21$\pm$0.83 35.25$\pm$0.46 36.17$\pm$0.91 33.24$\pm$1.24 32.51$\pm$0.65
+Ours 32.44$\pm$0.92 36.66$\pm$1.05 36.45$\pm$1.26 37.63$\pm$1.79 34.68$\pm$0.53 36.12$\pm$0.99 38.74$\pm$1.36 40.32$\pm$1.79
TopK 49.90$\pm$0.05 46.51$\pm$0.76 41.51$\pm$1.11 39.62$\pm$1.22 52.31$\pm$0.09 48.18$\pm$0.98 43.34$\pm$0.91 40.95$\pm$0.61
+Ours 49.75$\pm$0.41 49.62$\pm$0.62 46.48$\pm$1.07 44.44$\pm$1.40 51.73$\pm$0.49 51.19$\pm$1.08 48.21$\pm$1.01 45.52$\pm$1.40
DPP 50.56$\pm$0.06 46.59$\pm$0.79 42.54$\pm$1.09 40.26$\pm$1.22 53.66$\pm$0.15 48.38$\pm$0.98 43.53$\pm$0.91 41.83$\pm$0.64
+Ours 50.36$\pm$0.31 49.47$\pm$1.57 46.50$\pm$1.26 45.16$\pm$1.64 52.84$\pm$0.27 50.97$\pm$1.14 48.16$\pm$1.18 46.77$\pm$1.48
VoteK 48.69$\pm$0.14 45.71$\pm$0.88 41.51$\pm$1.13 39.94$\pm$0.95 52.62$\pm$0.24 48.02$\pm$0.71 42.72$\pm$0.91 40.57$\pm$0.60
+Ours 48.57$\pm$0.39 48.21$\pm$0.97 46.39$\pm$1.23 44.77$\pm$1.27 52.62$\pm$0.44 50.97$\pm$0.98 48.05$\pm$0.82 45.84$\pm$1.05
ConE 48.08$\pm$0.86 44.55$\pm$1.51 40.35$\pm$1.29 38.03$\pm$1.31 51.18$\pm$0.58 46.09$\pm$0.79 41.30$\pm$0.63 38.92$\pm$0.72
+Ours 48.51$\pm$0.70 46.89$\pm$0.85 45.09$\pm$1.08 43.21$\pm$1.32 50.95$\pm$0.46 48.32$\pm$0.84 46.83$\pm$1.07 44.13$\pm$0.79
ByDC 50.07$\pm$0.45 45.00$\pm$0.78 41.10$\pm$1.23 38.90$\pm$1.36 53.15$\pm$0.44 46.16$\pm$0.75 41.90$\pm$0.76 38.92$\pm$0.62
+Ours 49.82$\pm$0.39 48.54$\pm$0.94 46.60$\pm$1.11 44.52$\pm$1.05 52.61$\pm$0.55 49.98$\pm$1.00 47.33$\pm$1.23 44.39$\pm$1.14
Bayesian optimizing conditional bias
In practice, since the data distribution is unknown, we use empirical $\operatorname{Error}$ minimization of the balanced validation set $\mathcal{D}_{b}$ to achieve an empirical estimate of the underlying data distribution.
Typically, we minimize the empirical $\operatorname{Error}$ as follows:
$\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\{(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right].$
where we still use effective numbers (Cui et al., 2019) as class-wise weights where $w_{i}=(1-\alpha)/(1-\alpha^{n_{j}})$ and $\alpha=(N-1)/N$. We estimate the conditional bias $\beta$ by employing a Bayesian optimizing framework, which is suitable for non-differentiable and black-box function optimization. In this way, the estimated conditional bias $\beta$ can be regarded as an approximation of the conditional bias between annotated and test datasets. We present our overall optimizing process in Appendix A.3.
Re-weighting with adjusted class-wise weights
We use existing selection methods to select a subset of candidates (e.g. $K^{{}^{\prime}}$=100). We sort all candidates in increasing order by $(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ and obtain the indices for the sorted scores as $\mathcal{I}$. And we select these candidates with low-ranking examples in the sorted list $\mathcal{I}$:
$g(\mathrm{\mathbf{c}}_{i})=\mathrm{1}\left(\operatorname{Loc}(\mathrm{\mathbf{c}}_{i},\mathcal{I})\leq K\right),$ (8)
$\mathcal{I}=\operatorname{argsort}\left\{(w_{i}+\beta)s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t}))\right\}_{i=1}^{K^{{}^{\prime}}}.$ (9)

### v2

@@ 4 Methodology
Figure 2: The Kullback-Leibler (KL) divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class in the AgNews (a), Amazon (b) and Yelp (c) datasets. The KL divergence validates the existence of conditional bias in imbalanced datasets.
@@ 4.1 Decomposition of the distributional differences
For class-imbalanced annotated datasets $\mathcal{D}_{c}$, our goal is to select suitable demonstrations for a given test input $\mathrm{\mathbf{x}}_{t}$ in order to minimize the expected error $\operatorname{Error}$ across all classes:
@@
$\operatorname{Error}=\mathbb{E}_{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right],$ (7)
where $M[\cdot,\cdot]$ measures the mismatched degree between real output and the generated output conditioned on demonstrations $\mathrm{\mathbf{C}}_{K}$ and test input $\mathrm{\mathbf{x}}_{t}$. For example, $M[\cdot,\cdot]$ is the error rate for classification tasks and the negative value of the EM score for question-answering tasks.
Next, we apply the importance sampling trick to connect the expected $\operatorname{Error}$ of ICL with the annotated dataset to test dataset as follows:
@@
$\operatorname{Error}$ $=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\frac{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})},$
$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$ (8)
@@
where $P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})/P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})$ measures the distributional difference, $\bm{w}=P_{t}(Y)/P_{c}(Y)$ denotes class weights and $\bm{\beta}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ is conditional bias. The proof is shown in Appendix B.
For an imbalanced dataset, it is easily observed that there is a large discrepancy between the class priors of the annotated and test distributions, resulting in $\bm{w}\neq 1$. Due to target shift and concept shift caused by limited sampling in the tailed classes, the conditional distribution between the annotated and test datasets usually shows significant differences, thereby $\bm{\beta}\neq 0$. To verify the existence of conditional bias, we employ the Latent Dirichlet Allocation (LDA) model to derive the document-topic distributions for each class and then compute the Kullback-Leibler (KL) of conditional distribution between the annotated and test datasets. Figure 2 shows that the KL divergence of conditional distribution between the annotated and test datasets is not equal to zero, thereby verifying the existence of conditional bias $\bm{\beta}$ in the imbalanced datasets. Overall, the above analysis illustrates that the performance of ICL is affected by both class weights $\bm{w}$ and conditional bias $\bm{\beta}$; however, the latter is typically neglected in existing rebalancing methods in Subsection 3.2, leading to suboptimal performance of ICL with imbalanced annotations.
4.2 Reweighting with Conditional Bias
The previous analysis shows that the performance of ICL significantly deteriorates when demonstrations are selected from an imbalanced dataset without considering the conditional bias $\bm{\beta}$. However, directly computing the conditional bias $\bm{\beta}$ is non-trivial as real conditional distributions of annotated and test datasets are unknown. In what follows, we propose Reweighting with Conditional Bias (RCB), a complementary strategy to enhance the ICL performance under class imbalance. Our key idea is to incorporate conditional bias into the demonstration reweighting in selection. With this in mind, we present the details of our approach in the following.
Selecting a balanced subset
@@ 4.2 Reweighting with Conditional Bias
Given an imbalanced dataset $\mathcal{D}_{c}$ with $n_{j}$ examples for the $j$-th class, we select a balanced subset $\mathcal{D}_{b}$ by uniformly sampling $n_{b}$ examples from per class of $\mathcal{D}_{c}$:
$\mathcal{D}_{b}=\operatorname{UniformSample}(\mathcal{D}_{c},n_{b}),$ (9)
where $\operatorname{UniformSample}(\mathcal{D}_{c},n_{b})$ denotes uniformly selecting $n_{b}$ examples per class of $\mathcal{D}_{c}$, $n_{b}$ is constrained by the condition $n_{b}<\min_{j\in\mathcal{Y}}n_{j}$ to ensure $n_{b}$ does not exceed the size of the smallest class. Through the selection, we construct a balanced subset where all classes contribute an equal number of examples, aligning with the composition of the test dataset. It is worth noting that the selected balanced subset can be also used.
Estimating the conditional bias
We let the class weights $\bm{w}$ resemble the empirically successful design from previous literature [9, 28]. We continue to use effective numbers [9] as class weights $\bm{w}$, defined as $w_{j}=(1-\alpha)/(1-\alpha^{n_{j}})$ where $\alpha=(N-1)/N$ and $N$ denotes the total number of examples in imbalanced dataset $\mathcal{D}_{c}$. In practice, since the true conditional distribution cannot be directly computed, we approximate the conditional bias $\bm{\beta}$ by leveraging the balanced subset $\mathcal{D}_{b}$:
$\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\{(\bm{w}+\bm{\beta})\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right],$ (10)
where we compute the mismatched degree $M$ of the balanced subset $\mathcal{D}_{b}$ between real output $\bm{y}_{b}$ and ICL’s output conditioned on demonstrations selected from a remaining imbalanced dataset $\mathcal{D}_{r}=\mathcal{D}_{c}\setminus\mathcal{D}_{b}$, while preserving the original imbalance. We employ $\operatorname{Top}_{K}$ to select $K$ highest-ranked demonstrations based on reweighting existing demonstration scores for selection $s(\cdot,\cdot)$ via $\bm{w}+\bm{\beta}$. We use a Bayesian optimization framework, which is suitable for non-differentiable and black-box function optimization, to estimate conditional bias $\bm{\beta}$ with the lowest mismatched degree on the balanced subset. We show the detailed bayesian optimization process in Appendix C. In this way, we can estimate conditional bias $\bm{\beta}$ to measure the bias of feature distributions within classes.
Re-weighting with class weights $\bm{w}$ and conditional bias $\bm{\beta}$
Given a test input $\mathrm{\mathbf{x}}_{t}$, we first select $K^{\prime}$ candidates from the imbalanced dataset $\mathcal{D}_{c}$ using existing selection methods. A large $K^{\prime}$ ensures that all classes (especially tailed classes) are included in the subset of candidates selected. For each candidate $\mathbf{c}_{i}$, we use $\bm{w}+\bm{\beta}$ to reweight its original score $s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ and employ $\operatorname{argsort}$ to sort the adjusted score in ascending order, producing sorted indices $\mathcal{I}$:
$\mathcal{I}=\operatorname{argsort}\left\{\right(\bm{w}+\bm{\beta}\left)\times s\right(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t}\left)\right\}_{i=1}^{K^{{}^{\prime}}},$
We select these candidates with $K$ low-ranking indices in the sorted list $\mathcal{I}$ as the final demonstrations:
$g(\mathrm{\mathbf{c}}_{i})=\mathbbm{1}\left(\operatorname{Loc}(\mathrm{\mathbf{c}}_{i},\mathcal{I})\leq K\right).$
where $\mathbbm{1}$ is the indicator function and $\operatorname{Loc}(\mathrm{\mathbf{c}}_{i},\mathcal{I})$ return the index of $\mathrm{\mathbf{c}}_{i}$ in the sorted list $\mathcal{I}$. After the above steps, we establish the final $K$ demonstration set for in-context learning. Noticeably, our method offers several compelling advantages:
  • Algorithm-agnostic: Our method can be easily incorporated into existing demonstration selection methods, improving the performance of ICL with imbalanced annotations.
  • Easy to use: Our method requires access only to the model outputs and integrates effortlessly with any LLMs (see Figure 3 (d)). Our method is insensitive to the size of the balance subset $|\mathcal{D}_{b}|$ and class weights $\bm{w}$ (see Figure 3 (a) and (b)).
  • High efficiency: Our method only estimates the conditional bias for each class once for an imbalanced dataset. Thus, our method introduces negligible computational costs compared to in-context learning.
5 Experiments
5.1 Experimental Setup
Datasets and Evaluation For our evaluations, we verify the effectiveness of our method on four benchmark datasets, including Sentiment Classification (Amazon [10], Yelp [11]) and Topic Classification (AgNews, Yahoo [11]). Due to limited space, these tasks’ input/output, statistics and split are reported in the Appendix D. To simulate the issue of imbalanced data, we generate imbalanced datasets with pre-defined imbalance ratios $\phi$ (e.g., 1, 10, 50, 100) and we evaluate the performance of ICL on another corresponding class-balanced test dataset [24, 28]. We report average accuracy (%) with standard deviation (over 3 runs).
Models and Baselines We utilize various large language models, including open-weight models: OPT-6.7B, OPT-13B, OPT-30B, LLAMA-3-8B and LLAMA-3-70B; and APIs: ChatGPT-3.5-Turbo and Gemini-2.0-Flash. We use Bert-base-uncased sentence encoder [29] as the similarity tokenizer. We also conduct experiments with existing selection methods as baselines, including Random [30], TopK [4], DPP [17], VoteK [31], ConE [15], ByCS [6]. For hyperparameters, we set the size of the balanced subset as $|\mathcal{D}_{b}|=100$ and $K^{\prime}=1600$ by default. The details of our experiments are presented in Appendix D.
@@ 5.1 Experimental Setup
Imbalance Ratio 1 10 50 100 1 10 50 100
Random 47.70$\pm$0.77 43.72$\pm$1.26 39.59$\pm$0.92 38.01$\pm$0.58 50.68$\pm$0.49 46.65$\pm$0.72 41.30$\pm$0.99 39.90$\pm$0.54
+Ours 47.62$\pm$1.06 47.81$\pm$1.40 46.51$\pm$0.90 47.50$\pm$1.13 50.11$\pm$0.70 50.50$\pm$0.60 50.41$\pm$1.35 51.01$\pm$2.19
TopK 58.32$\pm$0.03 54.01$\pm$0.68 48.52$\pm$0.99 46.78$\pm$1.50 59.93$\pm$0.05 55.57$\pm$0.58 50.78$\pm$0.88 48.65$\pm$0.88
+Ours 58.12$\pm$0.65 57.46$\pm$0.50 54.14$\pm$1.31 51.92$\pm$1.22 59.53$\pm$0.37 58.94$\pm$0.93 55.58$\pm$0.75 52.83$\pm$1.32
DPP 58.50$\pm$0.00 53.90$\pm$0.79 49.37$\pm$0.78 47.25$\pm$1.11 60.00$\pm$0.07 55.54$\pm$0.85 51.30$\pm$0.74 49.58$\pm$0.58
+Ours 58.20$\pm$0.34 56.76$\pm$1.09 53.25$\pm$1.32 51.93$\pm$1.13 59.90$\pm$0.11 58.40$\pm$0.94 55.06$\pm$1.10 53.34$\pm$1.26
VoteK 58.07$\pm$0.14 53.51$\pm$0.78 48.75$\pm$0.92 47.21$\pm$1.26 60.00$\pm$0.15 55.53$\pm$0.49 50.46$\pm$0.79 48.42$\pm$0.90
+Ours 57.92$\pm$0.30 56.20$\pm$0.68 53.95$\pm$1.34 51.84$\pm$1.01 59.89$\pm$0.22 58.23$\pm$0.81 55.45$\pm$0.63 52.93$\pm$0.89
ConE 57.36$\pm$0.51 52.41$\pm$1.09 47.86$\pm$1.33 45.58$\pm$1.33 58.60$\pm$0.51 53.97$\pm$0.71 49.25$\pm$0.70 46.87$\pm$0.80
+Ours 57.59$\pm$0.62 54.75$\pm$0.91 52.45$\pm$0.70 50.12$\pm$1.29 58.85$\pm$0.41 56.06$\pm$0.81 54.19$\pm$0.74 51.17$\pm$0.75
ByDC 58.38$\pm$0.40 52.62$\pm$0.87 47.95$\pm$1.09 45.78$\pm$1.17 60.04$\pm$0.31 53.91$\pm$0.54 49.66$\pm$0.77 47.01$\pm$0.71
+Ours 58.16$\pm$0.37 57.24$\pm$0.78 53.25$\pm$1.27 51.45$\pm$0.93 60.03$\pm$0.28 57.53$\pm$0.89 54.53$\pm$0.70 51.44$\pm$0.79
Average 56.39$\pm$0.31 51.69$\pm$0.91 47.01$\pm$1.00 45.10$\pm$1.16 58.21$\pm$0.26 53.53$\pm$0.65 48.79$\pm$0.82 46.74$\pm$0.73
56.27$\pm$0.56 55.04$\pm$0.89 52.26$\pm$1.14 50.79$\pm$1.12 58.05$\pm$0.35 56.61$\pm$0.83 54.08$\pm$0.88 52.16$\pm$1.20
5.2 Main Results